OpenStudy (anonymous):

Malachi took cans to the recycling plant and put them into the can-crusher one at a time. After he had already crushed some of the cans, he started recording the number of cans he crushed per minute. The table shows how many cans Malachi put in the can-crusher at different times. What is the equation that represents Malachi’s situation? Record any fractional answers in decimal form. (y – 65) = (x – ) Malachi’s Situation Minutes 3 5 7 9 Cans crushed 49 65 81 97

OpenStudy (anonymous):

@triciaal

OpenStudy (anonymous):

Is the answer (y-65)=8(x-13)??

OpenStudy (anonymous):

@mLe

oregonduck (oregonduck):

differences in minutes: 2 2 2 (x differences) differences in cans crushed: 16 16 16 (y differences) Since the same differences in x causes same differences in y, this is a linear relationship. The slope of the straight line is: difference in y / difference in x = 16 / 2 = 8. The equation of the line in slope-point form is: y - y1 = m(x - x1) m = 8. If we choose (5,65) as the point from the table then: y - 65 = 8(x - 5)

OpenStudy (mle):

find the slope between the points first, can you do that?

OpenStudy (anonymous):

ok... i got it now i see the 5... simple error :/

OpenStudy (mle):

oh yeah, just put in the 5 and it'll be right ^_^

OpenStudy (anonymous):

lol... told ya... i have been at this since 6:30 am EST... running out of steam atm...lol

oregonduck (oregonduck):

lol

OpenStudy (triciaal):

sorry was away

OpenStudy (triciaal):

@OregonDuck thank you

oregonduck (oregonduck):

yup